Each value below represents a different aqueous solution at 25°C. Classify each solution as acidic, basic, or neutral. Acidicplease help me figure this out thanks!
Answer #1

Acidic Во с Neutral Си+] - 1:0 x 10-1 ри = 5. 20 ри = 13.98 рон = 4. 2.4 - 4 | Бон = 1. оо [н] = 5-7x 108 [н] : 1.8 x 10 [о